find the angles between the lines
A, x+2y=2 and 2x+4y=4
b,x+y=5 and x-y=4
c, x+2y=2 and 2x+y=3
2 find the distance between the lines
a,3x+2y=5 and -3x-2y=6
b, 5y=x and y=5x

Answer by Alan3354(69443)

find the angles between the lines
A, x+2y=2 and 2x+4y=4
That's the same line. The 2nd eqn is 2x the 1st.
-------------
b,x+y=5 and x-y=4
The 2 lines are parallel.
------------------------
c, x+2y=2 and 2x+y=3
x+2y=2 --> m = -1/2: arctan(-1/2) = -26.565º with the x-axis
2x+y=3 --> m = -2: arctan(-2) = -63.4349º with the x-axis
36.7º between the 2 lines
--------------------------
2 find the distance between the lines
a,3x+2y=5 and -3x-2y=6
b, 5y=x and y=5x